# Lower bound on the number of fixed points for circle actions on 10-dimensional almost complex manifolds

### v1

@@ 1. Introduction
Let the circle group act on a $2n$-dimensional compact almost complex manifold $M$. Let $k$ be the number of fixed points. When $k$ is small, possible dimensions for $M$ and known examples of such a manifold $M$ are as in Table 1; also see Theorem 2.8. Circle actions with few fixed points on different types of manifolds have been studied; for those results, we refer to [A, DW, GS, J1, J3, J4, J6, J8, KL, Kos1, Kos2, M, PT, T, W].
@@
Let $M$ be a compact almost complex manifold endowed with a circle action. If the number of fixed points is odd, then the dimension of the manifold is a multiple of 4; see Lemma 2.7. Now, suppose that the dimension of $M$ is 10. By the discussions above, $k$ is even and $k\neq 2$. In this paper, we establish an exact lower bound for the number of fixed points for a 10-dimensional manifold.
@@
Proof.
The examples of 10-dimensional $M$ with 6 fixed points are the standard linear action on $\mathbb{CP}^{5}$ and a diagonal action on $\mathbb{CP}^{2}\times S^{6}$; note that the Todd genus of $\mathbb{CP}^{5}$ is 1, while that of $\mathbb{CP}^{2}\times S^{6}$ is 0. Circle actions on complex manifolds and symplectic circle actions on symplectic manifolds are particular examples of circle actions on almost complex manifolds. Thus, the conclusion of Theorem 1.1 holds for complex manifolds and symplectic manifolds.
@@ 12 $S^{6}\times S^{6}$
The main ingredients for the proof of Theorem 1.2 are the Chern numbers, the Todd genus, and the ABBV localization theorem (Theorem 2.1). We outline the proof of Theorem 1.2. For this, let the circle group act on a 10-dimensional compact almost complex manifold $M$ with 4 fixed points. By using the previous work of [J5] and the ABBV localization theorem, we prove Lemma 3.4 that the Chern numbers $\int_{M}c_{1}^{3}c_{2}$ and $\int_{M}c_{1}^{2}c_{3}$ vanish. In Lemma 4.1, we show that three possible cases occur for the number $N_{i}$ of fixed points with $i$ negative weights. In each case, we use [GS, Theorem 1.2] that computes the Chern number $\int_{M}c_{1}c_{4}$ in terms of $N_{i}$. The Todd genus $\mathrm{Todd}(M)$ of $M$ satisfies $\mathrm{Todd}(M)=\int_{M}\frac{1}{1440}(-c_{1}c_{4}+c_{1}^{2}c_{3}+3c_{1}c_{2}^{2}-c_{1}^{3}c_{2})$. Using all of these, we determine the Chern number $\int_{M}c_{1}c_{2}^{2}$. In each possible case for $N_{i}$, we get a rational number for $\int_{M}c_{1}c_{2}^{2}$. This contradicts the fact that each Chern number is an integer. It follows that such a manifold $M$ does not exist.
@@ Acknowledgements
The author would like to thank Leonor Godinho and Silvia Sabatini for fruitful discussions and their warm hospitalities; this work was partially done during the author’s visits to Instituto Superior Técnico and University of Cologne.
@@ 2. Background and preliminaries
[AB] Let the circle act on a compact oriented manifold $M$. For any $\alpha\in H_{S^{1}}^{*}(M;\mathbb{Q})$,
@@
The below lemma is an easy consequence of Theorem 2.2.
Lemma 2.4.
[J7] Let a $k$-dimensional torus $T^{k}$ act on a $2n$-dimensional compact almost complex manifold $M$ with isolated fixed points. Let $\chi_{y}(M)=\sum_{i=0}^{n}a_{i}(M)\cdot(-y)^{i}$ be the Hirzebruch $\chi_{y}$-genus of $M$. Then the following hold:
  1. (1)
    $a_{i}(M)\geq 0$ for $0\leq i\leq n$.
  2. (2)
    The total number of fixed points is $\sum_{i=0}^{n}a_{i}(M)$.
  3. (3)
    For $0\leq i\leq n$, $a_{i}(M)=a_{n-i}(M)$.
@@
A classification of a circle action on a compact almost complex manifold with few fixed points is as follows.
@@
[J3] Let the circle act on a compact almost complex manifold $M$.
@@ 3. Chern numbers of high dimensional manifolds with 4 fixed points
Lemma 3.1 implies the following lemma.
@@
In this section, we prove Theorem 1.2. First, we consider possibilities for the Hirzebruch $\chi_{y}$-genus of a 10-dimensional compact almost complex $S^{1}$-manifold with 4 fixed points.
Lemma 4.1.
Let the circle act on a 10-dimensional compact almost complex manifold with exactly 4 fixed points. Then one of the following cases holds for the numbers $N_{i}$ of fixed points that have exactly $i$ negative weights.
  1. (1)
    $N_{0}=N_{1}=N_{4}=N_{5}=1$ and $N_{2}=N_{3}=0$.
  2. (2)
    $N_{0}=N_{5}=0$ and $N_{1}=N_{2}=N_{3}=N_{4}=1$.
  3. (3)
    $N_{0}=N_{1}=N_{4}=N_{5}=0$ and $N_{2}=N_{3}=2$.
Proof.
By definition, $N_{i}\geq 0$ for $0\leq i\leq 5$. Since there are exactly 4 fixed points, $\sum_{i=0}^{5}N_{i}=4$. By Theorem 2.2, $N_{0}=N_{5}$, $N_{1}=N_{4}$, and $N_{2}=N_{3}$.
First, suppose that $N_{0}>0$. Assume that $N_{0}\geq 2$. Since $N_{5}=N_{0}$, it follows that $N_{5}\geq 2$ and $N_{i}=0$ for $1\leq i\leq 4$. This contradicts Lemma 2.5 that there exists some $i$ such that $N_{i}\neq 0$ and $N_{i+1}\neq 0$. Therefore, $N_{0}=1$ and hence $N_{5}=N_{0}=1$. Together with $N_{i}\geq 0$ for $0\leq i\leq 5$ and $\sum_{i=0}^{5}N_{i}=4$, Lemma 2.5 implies that $N_{1}>0$ or $N_{4}>0$ and hence $N_{1}=N_{4}=1$ since $N_{1}=N_{4}$.
Second, suppose that $N_{0}=0$ and $N_{1}>0$. Then $N_{4}=N_{1}>0$. Lemma 2.5 implies that $N_{2}>0$ or $N_{3}>0$. Since $N_{2}=N_{3}$, it follows that $N_{1}=N_{2}=N_{3}=N_{4}=1$.
Third, suppose that $N_{0}=0$ and $N_{1}=0$. Since $\sum_{i=0}^{5}N_{i}=4$ and $N_{2}=N_{3}$, it follows that $N_{2}=N_{3}=2$.  $\Box$
With all of the above, we are ready to prove Theorem 1.2.
@@
Assume on the contrary that there exists a circle action on a 10-dimensional compact almost complex manifold $M$ exactly 4 fixed points. The Todd genus $\mathrm{Todd}(M)$ of $M$ is
@@
Moreover, since the Todd genus $\mathrm{Todd}(M)$ of $M$ is the Hirzebruch $\chi_{y}$-genus $\chi_{y}(M)$ of $M$ evaluated at $y=0$, by Theorem 2.2, $\mathrm{Todd}(M)=N_{0}$.
By Lemma 4.1, one of the following holds for $N_{i}$, where $N_{i}$ denotes the number of fixed points that have exactly $i$ negative weights.
  1. (1)
    $N_{0}=N_{1}=N_{4}=N_{5}=1$ and $N_{2}=N_{3}=0$.
  2. (2)
    $N_{0}=N_{5}=0$ and $N_{1}=N_{2}=N_{3}=N_{4}=1$.
  3. (3)
    $N_{0}=N_{1}=N_{4}=N_{5}=0$ and $N_{2}=N_{3}=2$.
First, suppose that Case (1) holds. In this case, $\mathrm{Todd}(M)=N_{0}=1$. By Theorem 2.3,
$\int_{M}c_{1}c_{4}=\sum_{i=0}^{5}N_{i}[6i(i-1)-25]=1\cdot[6\cdot 0\cdot(-1)-25]+1\cdot[6\cdot 1\cdot 0-25]+1\cdot[6\cdot 4\cdot 3-25]+1\cdot[6\cdot 5\cdot 4-25]=92.$
Therefore,
$1=\mathrm{Todd}(M)=\frac{1}{1440}(-92+\int_{M}3c_{1}c_{2}^{2})$
and thus $\int_{M}c_{1}c_{2}^{2}=\frac{1532}{3}$, which leads to a contradiction since any Chern number is an integer.
Second, suppose that Case (2) holds. In this case, $\mathrm{Todd}(M)=N_{0}=0$. By Theorem 2.3,
$\int_{M}c_{1}c_{4}=\sum_{i=0}^{5}N_{i}[6i(i-1)-25]=1\cdot[6\cdot 1\cdot 0-25]+1\cdot[6\cdot 2\cdot 1-25]+1\cdot[6\cdot 3\cdot 2-25]+1\cdot[6\cdot 4\cdot 3-25]=20.$
Therefore,
$0=\mathrm{Todd}(M)=\frac{1}{1440}(-20+\int_{M}3c_{1}c_{2}^{2})$
and thus $\int_{M}c_{1}c_{2}^{2}=\frac{20}{3}$, which leads to a contradiction.
Third, suppose that Case (3) holds. Then $\mathrm{Todd}(M)=N_{0}=0$. By Theorem 2.3,
$\int_{M}c_{1}c_{4}=\sum_{i=0}^{5}N_{i}[6i(i-1)-25]=2[6\cdot 2\cdot 1-25]+2[6\cdot 3\cdot 2-25]=-4.$
Therefore,
$0=\mathrm{Todd}(M)=\frac{1}{1440}(4+\int_{M}3c_{1}c_{2}^{2})$
and thus $\int_{M}c_{1}c_{2}^{2}=-\frac{4}{3}$, which leads to a contradiction. Therefore, this theorem holds.  $\Box$

### v2

@@ Abstract.
2020 Mathematics Subject Classifications: 58C30, 32Q60, 37C25
keywords: almost complex manifold, circle action, fixed point, Chern number
@@ 1. Introduction
Let the circle group act on a $2n$-dimensional compact, connected almost complex manifold $M$. Let $k$ be the number of fixed points. When $k$ is small, possible dimensions for $M$ and known examples of such a manifold $M$ are as in Table 1; also see Theorem 2.7. Circle actions with few fixed points on different types of manifolds have been studied; for those results, we refer to [A, DW, GS, J1, J3, J4, J6, J7, KL, Kos1, Kos2, LM, M, PT, T, W].
@@
Let $M$ be a compact, connected almost complex manifold endowed with a circle action. If the number $k$ of fixed points is odd, then the dimension of the manifold is a multiple of $4$; see Lemma 2.6. If $k=2$, by Theorem 2.7 (2), $M$ is the $2$-sphere or $\dim M=6$. Now, suppose that the dimension of $M$ is 10. By the discussions, $k$ is even and $k\neq 2$. In this paper, we establish an exact lower bound for the number of fixed points for a 10-dimensional manifold.
@@
We prove Theorem 1.1 in the last section. The examples of 10-dimensional $M$ with 6 fixed points are the standard linear action on $\mathbb{CP}^{5}$ and a diagonal action on $\mathbb{CP}^{2}\times S^{6}$; note that the Todd genus of $\mathbb{CP}^{5}$ is 1, while that of $\mathbb{CP}^{2}\times S^{6}$ is 0. Circle actions on complex manifolds and symplectic circle actions on symplectic manifolds are particular examples of circle actions on almost complex manifolds. Thus, the conclusion of Theorem 1.1 holds for complex manifolds and symplectic manifolds.
@@ 12 $S^{6}\times S^{6}$
The main ingredients for the proof of Theorem 1.2 are the Chern numbers, the Todd genus, and the ABBV localization theorem (Theorem 2.1). We outline the proof of Theorem 1.2. For this, let the circle group act on a $10$-dimensional compact almost complex manifold $M$ with 4 fixed points. The Todd genus $\mathrm{Todd}(M)$ of $M$ satisfies $\mathrm{Todd}(M)=\int_{M}\frac{1}{1440}(-c_{1}c_{4}+c_{1}^{2}c_{3}+3c_{1}c_{2}^{2}-c_{1}^{3}c_{2})$. By using the previous work of [J5] and the ABBV localization theorem, we prove Lemma 3.4 that the Chern numbers $\int_{M}c_{1}^{3}c_{2}$ and $\int_{M}c_{1}^{2}c_{3}$ vanish. Then we use [GS, Theorem 1.2], which computes the Chern number $\int_{M}c_{1}c_{4}$ in terms of the number $N_{i}$ of fixed points with $i$ negative weights, and get a rational number for $\int_{M}c_{1}c_{2}^{2}$. This contradicts the fact that each Chern number of a compact almost complex manifold is an integer. It follows that such a manifold $M$ does not exist.
@@ Acknowledgements
The author would like to thank Leonor Godinho and Silvia Sabatini for fruitful discussions and their warm hospitalities; this work was partially done during the author’s visits to Instituto Superior Técnico and University of Cologne. The author also thanks the anonymous referees for their valuable comments and suggestions, which helped improve the quality of the paper and led to a simpler proof of Theorem 1.2.
@@ 2. Background and preliminaries
[AB, BV] Let the circle act on a compact oriented manifold $M$. For any $\alpha\in H_{S^{1}}^{*}(M;\mathbb{Q})$,
@@
A classification of a circle action on a compact, connected almost complex manifold with few fixed points is as follows.
@@
[J3] Let the circle act on a compact, connected almost complex manifold $M$.
@@ 3. Chern numbers of high dimensional manifolds with 4 fixed points
In both cases of Lemma 3.1, the following holds.
@@
In this section, we prove Theorems 1.1 and Theorem 1.2. First, we prove Theorem 1.2.
@@
Assume on the contrary that there exists a circle action on a 10-dimensional compact almost complex manifold $M$ with exactly 4 fixed points. The Todd genus $\mathrm{Todd}(M)$ of $M$ is
@@
Moreover, since the Todd genus $\mathrm{Todd}(M)$ of $M$ is the Hirzebruch $\chi_{y}$-genus $\chi_{y}(M)$ of $M$ evaluated at $y=0$, by Theorem 2.2, $\mathrm{Todd}(M)=N_{0}$. Therefore, the following equation holds.
$\int_{M}3c_{1}c_{2}^{2}=1440\cdot N_{0}+\int_{M}c_{1}c_{4}$.
Since any Chern number of a compact almost complex manifold is an integer,
$\int_{M}3c_{1}c_{2}^{2}\equiv 0\pmod{3}$.
By Theorem 2.3, $\int_{M}c_{1}c_{4}=\sum_{i=0}^{5}N_{i}[6i(i-1)-25]$ and hence
$\int_{M}3c_{1}c_{2}^{2}=1440\cdot N_{0}+\int_{M}c_{1}c_{4}=1440\cdot N_{0}+\sum_{i=0}^{5}N_{i}\left[6i(i-1)-25\right]\equiv-\sum_{i=0}^{5}N_{i}=-4\pmod{3}$,
which is a contradiction. Therefore, this theorem holds.  $\Box$
Proof of Theorem 1.1.